OpenStudy (anonymous):

Write the equation that is the translation of y=|x| left 1 unit and up 2 units

OpenStudy (anonymous):

To translate a graph 2 units "up", you replace "y" with "y - 2" To translate a graph 1 unit to the left, you replace "x" with "x + 1" So, you would have: y - 2 = |x + 1| which can also be written: y = |x + 1| + 2 If you want to remove the absolute value sign, you can always write the function as 2 parts: y = x + 3 for x >= -1 and y = -x + 1 for x < -1

OpenStudy (anonymous):

@Friendsandfoes

OpenStudy (anonymous):

I think it would be y=|x+1|+2

OpenStudy (anonymous):

oh, awesome! i should be more selfassure of myself. Thanks Tcarroll. :)

OpenStudy (anonymous):

uw!